OpenStudy (anonymous):

implicity derive - 7x^3+6y^4 = 21

OpenStudy (anonymous):

\[7x ^{3}+6y ^{4} =21\]

OpenStudy (anonymous):

how would you type this in wolfram

OpenStudy (amistre64):

derive it as usual then solve for y'

OpenStudy (amistre64):

d/dx (f(x,y))

OpenStudy (anonymous):

= sign confusing me

OpenStudy (amistre64):

you dont derive an = sign :)

OpenStudy (anonymous):

haha i know just confusing me

OpenStudy (amistre64):

7x^3+6y^4 = 21 d(7x^3)/dx = ?

OpenStudy (anonymous):

21x^2 +24y^3 * dy/dx = 0 now make dy/dx the subject

OpenStudy (anonymous):

6y^4

OpenStudy (anonymous):

what about the 21

OpenStudy (amistre64):

we can get to that ... \[\frac{d(7x^3)}{dx} = ?\]

OpenStudy (anonymous):

7x^2

OpenStudy (anonymous):

the derivative of 21 is 0

OpenStudy (anonymous):

would i just write d/dx 21

myininaya (myininaya):

why do people say derive when they mean differentiate?

OpenStudy (amistre64):

not quite, and you are missing a part: \[\frac{d(7x^3)}{dx} = 21x^2\frac{dx}{dx}\]

OpenStudy (anonymous):

same thing

OpenStudy (amistre64):

what is this now: \[\frac{d(6y^4)}{dx}=?\]

OpenStudy (zarkon):

they are different

OpenStudy (anonymous):

confused

myininaya (myininaya):

it drives me crazy a little lol

OpenStudy (anonymous):

minds are well tuned then =] how to solve this???

OpenStudy (amistre64):

\[\frac{d(6y^4)}{dx}=?\]

OpenStudy (anonymous):

mistre confused here

OpenStudy (amistre64):

youll have to be more specific, I can quite read your mind to see what you are actually confused about ...

OpenStudy (anonymous):

not sure wether its just d6y4/dx

OpenStudy (amistre64):

would it be more solvable for you if it were like this? \[\frac{d(6x^4)}{dx}=?\]

OpenStudy (anonymous):

not sure how to multiply that out

OpenStudy (zarkon):

how would you write \[\frac{d}{dx}6[f(x)]^4\]

myininaya (myininaya):

if you have h(x)=6(f(x))^4 and i asked you to find h' what would you say? hopefully you know to use chain rule and say h'(x)=6*4*(f(x))^3*f'(x)=24f(x)[f(x)]^3

myininaya (myininaya):

lol zarkon thats what i was gonna say or did

OpenStudy (anonymous):

yes now its just d5x^4

OpenStudy (zarkon):

nice ;)

myininaya (myininaya):

oops i missed my little prime thingy

OpenStudy (zarkon):

only on the second part ;)

OpenStudy (amistre64):

the rules for differentation dont change becuase the variable changes; 6x^4 derives in the exact same manner as 6y^4 right?

myininaya (myininaya):

h'(x)=24f'(x)[f(x)]^3

OpenStudy (anonymous):

thats true but dx/dy in implicit differentiation confuses me

OpenStudy (anonymous):

7x^3+6y^4=21 how to solve this one i have 2 more to post

OpenStudy (amistre64):

consider this then: \[\frac{d}{dx}(6x^4)=\frac{dx}{dx}(24x^3)\] out poped the variable, do you see it? happens all the time; but the thing is dx/dx = 1 so they toss it out. \[\frac{d}{dx}(6y^4)=\frac{dy}{dx}(24y^3)\] the variable still pops out of it, but we have to keep this derived bit left in the equation since it has no "real" value to associate it to

OpenStudy (anonymous):

ok understood that

OpenStudy (amistre64):

good :) and the 21 on the other side is just a constant; and all constant derive to 0

OpenStudy (anonymous):

so now how to write the whole thing out ?

OpenStudy (amistre64):

7x^3 +6y^4 = 21 ; derives to \[21x^2\frac{dx}{dx} + 24y^3 \frac{dy}{dx} = 0\] \[21x^2 + 24y^3 \frac{dy}{dx} = 0\] now we solve for dy/dx

OpenStudy (anonymous):

wouldnt it be 21x^3 or am i forgetting dx/dx

OpenStudy (amistre64):

- 21x^2 and divide out the 24y^3 right? \[\frac{dy}{dx}=\frac{-21x^2}{24y^3}\]

OpenStudy (anonymous):

right

OpenStudy (amistre64):

7x^3 just derives to: (3) 7x^2 (dx/dx) = 21x^2

OpenStudy (anonymous):

wouldnt 24 be on top half of the equation for that example

OpenStudy (amistre64):

no, the rest of it is just algebra .... ax + by = 0 by = -ax y = -ax/b

OpenStudy (anonymous):

right sorry about that

OpenStudy (anonymous):

will start a new thread